# Enhancing LLM-based Search Agents via Contribution Weighted Group Relative Policy Optimization

### 15 Apr 2026 v1

@@ Abstract
Search agents extend Large Language Models (LLMs) beyond static parametric knowledge by enabling access to up-to-date and long-tail information unavailable during pretraining. While reinforcement learning has been widely adopted for training such agents, existing approaches face key limitations: process supervision often suffers from unstable value estimation, whereas outcome supervision struggles with credit assignment due to sparse, trajectory-level rewards. To bridge this gap, we propose Contribution-Weighted GRPO (CW-GRPO), a framework that integrates process supervision into group relative policy optimization. Instead of directly optimizing process rewards, CW-GRPO employs an LLM judge to assess the retrieval utility and reasoning correctness at each search round, producing per-round contribution scores. These scores are used to rescale outcome-based advantages along the trajectory, enabling fine-grained credit assignment without sacrificing optimization stability. Experiments on multiple knowledge-intensive benchmarks show that CW-GRPO outperforms standard GRPO by 5.0% on Qwen3-8B and 6.3% on Qwen3-1.7B, leading to more effective search behaviors. Additional analysis reveals that successful trajectories exhibit concentrated contributions across rounds, providing empirical insight into search agent tasks.
@@ 1 Introduction
Search agents empower large language models (LLMs) to move beyond fixed parametric knowledge by providing access to real-time information and specialized facts absent from pre-training [li2025searcho1agenticsearchenhancedlarge, zhang2025deep]. Through iterative retrieval of external evidence and integration into the reasoning process, search agents ground model predictions in verifiable sources and substantially enhance the factual reliability of LLM-based systems [zhao2026retrieval, li2025matching, zhou2024trustworthiness, wang2025actingreasoningmoreteaching]. Consequently, they have become a key role in enabling LLMs to perform knowledge-intensive and fact-sensitive tasks [glass2022re2g, sawarkar2024blended].
While reinforcement learning (RL) offers a significant paradigm for training such agents, current RL approaches suffer from fundamental limitations [lee2024multi]. Depending on the level of supervision, they can be broadly categorized into process supervision and outcome supervision [uesato2022solving]. Process supervision [lightman2023letsverifystepstep, zhu2025chain, wang2025stepsearchignitingllmssearch] assigns round-level rewards and is commonly optimized with actor–critic methods such as Proximal Policy Optimization [schulman2017proximalpolicyoptimizationalgorithms], but learning reliable critics over diverse intermediate states is unstable and frequently leads to biased advantages and brittle training [gurumurthy2023practical, hu2025reinforcestabilizingcriticfreepolicy]. Outcome supervision, by contrast, relies solely on final answer correctness, which makes it difficult to attribute task success to individual search rounds, leading to a credit assignment problem and obscuring the uneven contributions of intermediate decisions. Recent advances, such as Group Relative Policy Optimization [shao2024deepseekmathpushinglimitsmathematical], improve training stability and memory efficiency, making outcome-supervised optimization more practical; however, they[jin2025searchr1trainingllmsreason, li2025searcho1agenticsearchenhancedlarge, wang2025actingreasoningmoreteaching, li2025r3] do not alter the sparsity of the reward signal, and the credit assignment issue remains unresolved.
@@ 2 Related Work
2.0.0.1 Agentic Search.
Large language models encode knowledge in static parameters, limiting access to up-to-date and long-tail knowledge. Retrieval-augmented generation (RAG) addresses this by enabling models to access external knowledge sources [zhang2026opennoveltyllmpoweredagenticverifiable, zhao2026retrieval, li2025matching, zhou2024trustworthiness]. An early paradigm is search-as-a-tool, where LLMs are prompted or fine-tuned to interact with static corpora. IRCoT [trivedi2023interleavingretrievalchainofthoughtreasoning] and FLARE [zhang2025flarefeedforwardgeometryappearance] use prompts to guide iterative reasoning and relieve model uncertainty, and Self-RAG [asai2023selfraglearningretrievegenerate] adopts supervised fine-tuning with reflection tokens for adaptive retrieval. More recently, agentic search enables LLMs to interact with real-world search engines for dynamic information access. WebGPT [nakano2022webgptbrowserassistedquestionansweringhuman] pioneers this direction using RLHF. Search-o1 [li2025searcho1agenticsearchenhancedlarge] integrates agentic search into large reasoning models via prompting, while Search-R1 [jin2025searchr1trainingllmsreason] and R1-Searcher [song2025r1searcherincentivizingsearchcapability] adopt RL for end-to-end training with outcome-based rewards. However, outcome supervision assigns uniform credit across all search rounds, failing to distinguish pivotal searches from redundant ones.
2.0.0.2 Outcome Supervision and Process Supervision.
A key challenge in training search agents lies in credit assignment across multi-round search trajectories. Outcome supervision, widely adopted in RL methods like PPO [schulman2017proximalpolicyoptimizationalgorithms] and GRPO [shao2024deepseekmathpushinglimitsmathematical], assigns rewards only based on the final answer, treating all immediate rounds equally and ignoring their unequal contributions. Process supervision addresses this via round-level feedback for finer-grained credit assignment. Early approaches to process supervision rely on learned Process Reward Models (PRMs), but often suffer from costly round-level annotation requirements and limited out-of-domain generalizability [lightman2023letsverifystepstep, huang-etal-2025-empirical]. To avoid training explicit reward models, recent works increasingly leverage frozen LLMs as external evaluators to assess intermediate rounds. These methods use LLM-based judges to assess reasoning quality and then assign round-level rewards, enabling process supervision without training dedicated PRMs [zhang2025criticsearchfinegrainedcreditassignment]. These approaches typically optimize judge-produced signals directly as rewards, making them sensitive to evaluation noise and calibration errors. In contrast, CW-GRPO uses LLM-derived signals only to redistribute outcome-based advantages, avoiding direct optimization over noisy intermediate rewards and thereby enabling more stable training.
@@ 3.3.1 Signal Definition and Rubrics
The conjunctive contribution signal is defined as the logical product $p_{i}^{t}=u_{i}^{t}\cdot v_{i}^{t}$. By enforcing a binary bottleneck, we treat each search round as a discrete “gate”—either it contributes a necessary building block to the solution, or it does not.
@@ 3.3.2 Judge Calibration and Reliability
The rubrics were iteratively refined to minimize ambiguity. In our final setup, the LLM judge achieved a 95% consensus rate with human expert annotations on both retrieval utility and reasoning correctness. This high level of agreement provides a reliable foundation for the subsequent advantage redistribution, ensuring that policy gradients are directed by high-fidelity process signals.
@@ 3.4 Adaptive Weighting Mechanism
For successful trajectories ($R_{i}=1$), rounds that introduce new task-relevant evidence and maintain logically consistent reasoning are typically aligned with the final successful outcome. This enables reliable credit redistribution across rounds. A temperature-controlled softmax is therefore adopted to emphasize high-contribution rounds:
@@
For failed trajectories ($R_{i}=0$), round-level attribution is substantially more ambiguous. Many intermediate rounds exhibit reasonable behavior, such as attempting to retrieve missing evidence or issuing follow-up queries for verification. However, these rounds may still fail to obtain useful information because the required knowledge is not covered by the corpus or cannot be retrieved due to limitations of the retriever. In such cases, the failure does not reflect an identifiable error in the agent’s decision at a specific round. As a result, assigning differentiated weights based on round-level signals is unreliable and can introduce spurious supervision. A more detailed analysis of such failure patterns is provided in Appendix A. To mitigate this issue, we assign a uniform contribution across all rounds:
@@ 3.5 Advantage Reallocation and Optimization
The scaling factor $(T_{i}-1)$ ensures the magnitude of the total learning signal for the trajectory remains conserved, i.e.,
@@
In summary, CW-GRPO achieves a seamless integration of process supervision into the GRPO framework. Our method achieves effective credit assignment across iterative search rounds by reallocating trajectory-level advantages based on each round’s contribution, rather than estimating absolute intermediate rewards. This design preserves the training stability of group-relative optimization while ensuring the policy learns more effectively from the specific decisions that drive task success.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate our method on two categories of knowledge-intensive tasks: (1) General Question Answering: Natural Questions (NQ) [kwiatkowski-etal-2019-natural], TriviaQA [joshi2017triviaqalargescaledistantly], and PopQA [mallen2023trustlanguagemodelsinvestigating]. (2) Multi-Hop Question Answering: HotpotQA [yang2018hotpotqadatasetdiverseexplainable], 2WikiMultiHopQA [ho2020constructingmultihopqadataset], Musique [trivedi2022musiquemultihopquestionssinglehop], and Bamboogle [press2023measuring]. We merge the training sets of NQ and HotpotQA as the training data. To mitigate the impact of decoding variance, we use Avg@4 Exact Match (EM) as the evaluation metric, averaging the EM scores over four sampled responses per task.
4.1.0.1 Hard-Case Evaluation Set.
In standard settings, models often bypass external search by leveraging internal parametric knowledge, which obscures their true agentic search proficiency. In light of this, we adopt a more challenging protocol using AgentGym-SearchQA-test, a 400-sample test set from AgentGym-RL [xi2025agentgymrltrainingllmagents]. These samples are specifically filtered from cases where the Qwen2.5-72B-Instruct [yang2024qwen25mathtechnicalreportmathematical] model failed, representing the "hardest" distribution for current LLMs. This setup shifts the focus from simple parametric recall to complex information-seeking and reasoning capabilities, explicitly emphasizing whether agents retrieve and utilize external evidence rather than relying on parametric knowledge.
4.1.0.2 Strict Retrieval Constraint.
To further emphasize the model’s search agentic capabilities, we include a specific instruction in the system prompt: the model is prohibited from using any parametric knowledge beyond basic common sense.
@@ 4.2 Baselines
We compare our approach against two categories of search agents: (1) Outcome-Supervised Methods: Search-R1-PPO and Search-R1-GRPO [jin2025searchr1trainingllmsreason], which optimize the final answer via group relative policy optimization. (2) Process-Supervised Methods: R3-RAG [li2025r3] and MT-PPO [wei2025reinforcingmultiturnreasoningllm], which utilize explicit process rewards and PPO frameworks.
We also include results from state-of-the-art closed-source models (GPT-5.1 [singh2025openaigpt5card], Gemini-3-Pro-Preview [google2025gemini3pro], Qwen3-Max [qwen3max]) as strong upper-bound references, alongside powerful open-source models (DeepSeek-V3.2 [deepseekai2025deepseekv32pushingfrontieropen], Qwen3-32B [qwen3technicalreport]) for comparison.
4.2.0.1 Note on Comparison:
Because of our uniquely challenging test set and the strict use of a no-parametric-knowledge prompt, the baseline performance reported in this paper differs from that reported in the original publications.
@@ 4.3 Implementation Details
We use Qwen3-8B and Qwen3-1.7B [qwen3technicalreport] as base models due to their tool-calling capabilities. Training is conducted using the veRL [sheng2024hybridflow] framework. The maximum context length is set to 9192 tokens, with up to 10 interaction rounds. We employ SGLang [zheng2024sglang] as the inference engine for rollout, setting both temperature and top_p both to 1.0.
@@
4.3.0.1 Retrieval Infrastructure.
Following Search-R1, we use the 2018 Wikipedia dump [wikidump] as the knowledge source. We employ E5 [wang2022text] as the retriever, with the number of retrieved documents $k$ set to 3.
4.3.0.2 Training Hyperparameters.
For all RL-based experiments, we use a KL-divergence regularization coefficient $\beta=0.001$ and a clipping ratio $\epsilon=0.2$. Specifically, for our proposed CW-GRPO, we apply GPT-oss-120B [openai2025gptoss120bgptoss20bmodel] as the judge model and set $\alpha=\infty$. This configuration ensures that for trajectories resulting in task success, the learning signals during search iterations are exclusively concentrated on those rounds where the conjunctive contribution signal is $1$. Other specific parameters are detailed in Table 1. The differences in hyperparameter configurations between GRPO-based and PPO-based methods stem from their distinct algorithmic characteristics (e.g., GRPO relies on intra-group relative comparisons while PPO optimizes independent trajectories), and we ensure fairness by normalizing the total number of sampled trajectories per training step.
@@ 5 Experimental Results
Search-R1-PPO 19.5 50.5 25.0 25.5 24.5 5.0 14.0 21.06
Search-R1-GRPO 18.0 45.5 25.0 27.0 22.5 6.75 16.5 21.00
R3-RAG 20.0 47.5 25.5 22.0 18.5 5.25 13.5 19.69
MT-PPO 25.5 49.0 25.0 24.0 31.5 6.0 13.0 22.19
@@ 5.1 Main Results
5.1.0.1 Overall Performance.
Our experimental results demonstrate the robust superiority of CW-GRPO, which achieves state-of-the-art performance across both model scales, outperforming all outcome- and process-supervised baselines on the identical backbone. On Qwen3-8B, CW-GRPO achieves an overall score of 31.38, representing a $5.0\%$ relative improvement over Search-R1-GRPO (29.88). On Qwen3-1.7B, CW-GRPO improves the overall performance from 21.00 to 22.31, corresponding to a $6.3\%$ relative gain. These results indicate that CW-GRPO scales favorably across model sizes and is particularly effective in low-capacity regimes, where efficient search credit assignment is critical.
5.1.0.2 Comparison with Open- and Closed-Source Models.
With Qwen3-8B as the backbone, CW-GRPO surpasses all evaluated open-source models, including DeepSeek-V3.2-Thinking and Qwen3-32B, despite operating under a stricter evaluation setting. However, a clear performance gap remains between CW-GRPO and large closed-source models such as GPT-5.1, Gemini-3-Pro-Preview, and Qwen3-Max, suggesting that stronger base models and proprietary training data still provide advantages under this challenging hard-case distribution.
5.1.0.3 General QA vs. Multi-Hop QA.
CW-GRPO demonstrates consistent improvements on both General QA and Multi-Hop QA tasks compared to outcome-supervised and process-supervised baselines. On Multi-Hop QA, CW-GRPO yields the most pronounced gains. On both Qwen3-1.7B and Qwen3-8B, it improves performance on HotpotQA, Musique, and Bamboogle simultaneously, indicating stronger long-horizon reasoning and evidence aggregation. On General QA, CW-GRPO remains competitive across datasets. While it achieves the best overall General QA performance on Qwen3-8B, its improvement on Qwen3-1.7B is smaller than that of MT-PPO. This suggests that for simpler single-hop questions, dense process supervision can still be effective for very small models, whereas CW-GRPO shows clearer advantages as reasoning depth increases.
5.1.0.4 Summary.
Overall, these results demonstrate that CW-GRPO provides a robust and scalable improvement over existing outcome- and process-supervised search agents, particularly for multi-hop reasoning under strict retrieval constraints. The consistent gains across model sizes highlight the effectiveness of contribution-weighted optimization in guiding search behavior, even in extremely challenging evaluation settings.

### 20 Apr 2026 v2

@@ Abstract
Search agents extend Large Language Models (LLMs) beyond static parametric knowledge by enabling access to up-to-date and long-tail information unavailable during pretraining. While reinforcement learning has been widely adopted for training such agents, existing approaches face key limitations: process supervision often suffers from unstable value estimation, whereas outcome supervision struggles with credit assignment due to sparse, trajectory-level rewards. To bridge this gap, we propose Contribution-Weighted GRPO (CW-GRPO), a framework that integrates process supervision into group relative policy optimization. Instead of directly optimizing process rewards, CW-GRPO employs an LLM judge to assess the retrieval utility and reasoning correctness at each search round, producing per-round contribution scores. These scores are used to rescale outcome-based advantages along the trajectory, enabling fine-grained credit assignment without sacrificing optimization stability. Experiments on multiple knowledge-intensive benchmarks show that CW-GRPO outperforms standard GRPO by 5.0% on Qwen3-8B and 6.3% on Qwen3-1.7B, leading to more effective search behaviors. Additional analysis reveals that successful trajectories exhibit concentrated contributions in specific rounds, providing empirical insight into search agent tasks.
@@ 1 Introduction
Search agents empower large language models (LLMs) to move beyond fixed parametric knowledge by providing access to real-time information and specialized facts absent from pre-training [Li et al., 2025a, Zhang et al., 2025a]. Through iterative retrieval of external evidence and integration into the reasoning process, search agents ground model predictions in verifiable sources and substantially enhance the factual reliability of LLM-based systems [Zhao et al., 2026, Li et al., 2025b, Zhou et al., 2024, Wang et al., 2025a]. Consequently, they have become a key role in enabling LLMs to perform knowledge-intensive and fact-sensitive tasks [Glass et al., 2022, Sawarkar et al., 2024].
While reinforcement learning (RL) offers a significant paradigm for training such agents, current RL approaches suffer from fundamental limitations [Jin et al., 2025a]. Depending on the level of supervision, they can be broadly categorized into process supervision and outcome supervision [Uesato et al., 2022]. Process supervision [Lightman et al., 2023, Zhu et al., 2025, Wang et al., 2025b] assigns round-level rewards and is commonly optimized with actor–critic methods such as Proximal Policy Optimization [Schulman et al., 2017], but learning reliable critics over diverse intermediate states is unstable and frequently leads to biased advantages and brittle training [Liu et al., 2024, Kazemnejad et al., 2025]. Outcome supervision, by contrast, relies solely on final answer correctness, which makes it difficult to attribute task success to individual search rounds, leading to a credit assignment problem and obscuring the uneven contributions of intermediate decisions. Recent advances, such as Group Relative Policy Optimization [Shao et al., 2024], improve training stability and memory efficiency, making outcome-supervised optimization more practical; however, they [Jin et al., 2025b, Wang et al., 2025a] do not alter the sparsity of the reward signal, and the credit assignment issue remains unresolved.
@@ 2 Related Work
Agentic Search. Large language models encode knowledge in static parameters, limiting access to up-to-date and long-tail knowledge. Retrieval-augmented generation (RAG) addresses this by enabling models to access external knowledge sources [Zhang et al., 2026, Zhao et al., 2026, Li et al., 2025b, Zhou et al., 2024]. Pioneering retrieval-augmented reasoning works such as IRCoT [Trivedi et al., 2023] and FLARE [Jiang et al., 2023] use prompts to guide iterative reasoning and reduce model uncertainty, and Self-RAG [Asai et al., 2024] adopts supervised fine-tuning with reflection tokens for adaptive retrieval. More recently, agentic search enables LLMs to interact with real-world search engines for dynamic information access. WebGPT [Nakano et al., 2022] pioneers this direction using RLHF. Search-o1 [Li et al., 2025a] integrates agentic search into large reasoning models via prompting, while Search-R1 [Jin et al., 2025b], R1-Searcher [Song et al., 2025] and MM-Doc-R1 [Lin et al., 2026] adopt RL for end-to-end training with outcome-based rewards. However, outcome supervision assigns uniform credit across search rounds, failing to distinguish pivotal searches from redundant ones.
Outcome Supervision and Process Supervision. A key challenge in training search agents lies in credit assignment across multi-round search trajectories. Outcome supervision, widely adopted in RL methods like PPO [Schulman et al., 2017] and GRPO [Shao et al., 2024], assigns rewards only based on the final answer, treating all immediate rounds equally and ignoring their unequal contributions. Process supervision addresses this via round-level feedback for finer-grained credit assignment. Early approaches to process supervision rely on learned Process Reward Models (PRMs), but often suffer from costly round-level annotation requirements and limited out-of-domain generalizability [Lightman et al., 2023, Huang et al., 2025]. To avoid training explicit reward models, recent works increasingly leverage frozen LLMs as external evaluators to assess intermediate rounds. These methods use LLM-based judges to assess reasoning quality and then assign round-level rewards, enabling process supervision without training dedicated PRMs [Zhang et al., 2025b]. These approaches typically optimize judge-produced signals directly as rewards, making them sensitive to evaluation noise and calibration errors. In contrast, CW-GRPO uses LLM-derived signals only to redistribute outcome-based advantages, avoiding direct optimization over noisy intermediate rewards and thereby enabling more stable training.
@@ 3.3.1 Signal Definition and Rubrics
The conjunctive contribution signal is defined as the logical product $p_{i}^{t}=u_{i}^{t}\cdot v_{i}^{t}$. By enforcing a binary bottleneck, we apply a discrete gate to each search round: it either contributes a necessary building block to the solution or it does not.
@@ 3.3.2 Judge Calibration and Reliability
The rubrics were iteratively refined to minimize ambiguity. In our final setup, the LLM judge achieved a 95% consensus rate with human expert annotations on both retrieval utility and reasoning correctness. This high level of agreement provides a reliable foundation for the subsequent advantage redistribution, ensuring that policy gradients are directed by high-fidelity process signals. Additional calibration details are deferred to Appendix B.
@@ 3.4 Adaptive Weighting Mechanism
For successful trajectories ($R_{i}=1$), rounds that introduce new task-relevant evidence and maintain logically consistent reasoning are typically associated with the final successful outcome. A temperature-controlled softmax is therefore adopted to emphasize high-contribution rounds:
@@
For failed trajectories ($R_{i}=0$), round-level attribution is substantially more ambiguous. Many intermediate rounds exhibit reasonable behavior, such as attempting to retrieve missing evidence or issuing follow-up queries for verification. However, these rounds may still fail to obtain useful information because the required knowledge is not covered by the corpus or cannot be retrieved due to limitations of the retriever. In such cases, the failure cannot be readily attributed to identifiable errors in the agent’s decisions at specific rounds. As a result, assigning differentiated weights based on round-level signals is unreliable and can introduce spurious supervision. A more detailed analysis of such failure patterns is provided in Appendix A. To mitigate this issue, we assign a uniform contribution across all rounds:
@@ 3.5 Advantage Reallocation and Optimization
The factor $(T_{i}-1)$ keeps the total learning signal magnitude of the trajectory unchanged, i.e.,
@@
In summary, CW-GRPO seamlessly integrates process supervision into the GRPO framework and achieves effective credit assignment across iterative search rounds by reallocating trajectory-level advantages based on each round’s contribution, rather than estimating absolute intermediate rewards. This design preserves the training stability of group-relative optimization while ensuring the policy learns more effectively from the specific decisions that drive task success.
@@ 4.1 Datasets and Evaluation Metrics
We evaluate our method on two categories of knowledge-intensive tasks: (1) General Question Answering: Natural Questions (NQ) [Kwiatkowski et al., 2019], TriviaQA [Joshi et al., 2017], and PopQA [Mallen et al., 2023]. (2) Multi-Hop Question Answering: HotpotQA [Yang et al., 2018], 2WikiMultiHopQA [Ho et al., 2020], Musique [Trivedi et al., 2022], and Bamboogle [Press et al., 2023]. We merge the training sets of NQ and HotpotQA as the training data. To mitigate the impact of decoding variance, we use Avg@4 Exact Match (EM) as the evaluation metric, averaging the EM scores over four sampled responses per task.
Hard-Case Evaluation Set. In standard settings, models often bypass external search by leveraging internal parametric knowledge, which obscures their true agentic search proficiency. In light of this, we adopt a more challenging protocol using AgentGym-SearchQA-test, a 400-sample test set from AgentGym-RL [Xi et al., 2025]. These samples are specifically filtered from cases where the Qwen2.5-72B-Instruct [Team, 2024] model failed, representing the "hardest" distribution for current LLMs. This setup shifts the focus from simple parametric recall to complex information-seeking and reasoning capabilities, explicitly emphasizing whether agents retrieve and utilize external evidence rather than relying on parametric knowledge.
Strict Retrieval Constraint. To further emphasize the model’s search agentic capabilities, we include a specific instruction in the system prompt: the model is prohibited from using any parametric knowledge beyond basic common sense.
@@ 4.2 Baselines
We compare our approach against two categories of search agents: (1) Outcome-Supervised Methods: Search-R1-PPO and Search-R1-GRPO [Jin et al., 2025b], which optimize the final answer via PPO and GRPO respectively. (2) Process-Supervised Methods: R3-RAG [Li et al., 2025c] and MT-PPO [Wei et al., 2025], which utilize explicit process rewards and PPO frameworks.
We also include results from state-of-the-art closed-source models (GPT-5.1 [Singh et al., 2025], Gemini-3-Pro-Preview [Google, 2025], Qwen3-Max [Team, 2025a]) as strong upper-bound references, alongside powerful open-source models (DeepSeek-V3.2 [DeepSeek-AI et al., 2025], Qwen3-32B [Team, 2025b]) for comparison.
Note on Comparison: Because of the uniquely challenging test set and the strict use of a no-parametric-knowledge prompt, the baseline performance reported in this paper differs from that reported in the original publications.
@@ 4.3 Implementation Details
We use Qwen3-8B and Qwen3-1.7B [Team, 2025b] as base models due to their tool-calling capabilities. Training is conducted using the veRL [Sheng et al., 2024] framework. The maximum context length is set to 9192 tokens, with up to 10 interaction rounds. We employ SGLang [Zheng et al., 2024] as the inference engine for rollout, setting both temperature and top_p to 1.0.
@@
Retrieval Infrastructure. We use the wiki-18-corpus dataset released by Search-R1 [Jin et al., 2025b], derived from the 2018 Wikipedia dump, as the knowledge source, and E5-base-v2 [Wang et al., 2022] as the retriever, with the number of retrieved documents $k$ set to 3.
Training Hyperparameters. For all RL-based experiments, we use a KL-divergence regularization coefficient $\beta=0.001$ and a clipping ratio $\epsilon=0.2$. Specifically, for our proposed CW-GRPO, we apply GPT-oss-120B [Agarwal et al., 2025] as the judge model and set $\alpha=\infty$. This configuration ensures that for trajectories resulting in task success, the learning signals during search iterations are exclusively concentrated on those rounds where the conjunctive contribution signal is $1$. Other specific parameters are detailed in Table 1. The hyperparameter differences between GRPO-based and PPO-based methods stem from their distinct algorithmic characteristics (intra-group relative comparisons vs. independent trajectory optimization), and we ensure fairness by normalizing the total number of sampled trajectories per training step.
@@ 5 Experimental Results
Search-R1-PPO 19.5 50.5 25.0 25.5 24.0 5.0 14.0 21.06
Search-R1-GRPO 18.0 45.5 25.0 27.0 22.5 6.8 16.5 21.00
R3-RAG 20.0 47.5 25.5 22.0 18.5 5.3 13.5 19.69
MT-PPO 22.5 49.0 25.0 24.0 31.5 6.0 13.5 22.19
@@ 5.1 Main Results
Overall Performance. Our experimental results demonstrate the robust superiority of CW-GRPO, which achieves state-of-the-art performance across both model scales, outperforming all outcome- and process-supervised baselines on the identical backbone. On Qwen3-8B, CW-GRPO achieves an overall score of 31.38, representing a $5.0\%$ relative improvement over Search-R1-GRPO (29.88). On Qwen3-1.7B, CW-GRPO improves the overall performance from 21.00 to 22.31, corresponding to a $6.3\%$ relative gain. These results indicate that CW-GRPO scales favorably across model sizes and is particularly effective in low-capacity regimes, where efficient search credit assignment is critical.
Comparison with Open- and Closed-Source Models. With Qwen3-8B as the backbone, CW-GRPO surpasses all evaluated open-source models, including DeepSeek-V3.2-Thinking and Qwen3-32B, despite operating under a stricter evaluation setting. However, a clear performance gap remains between CW-GRPO and large closed-source models such as GPT-5.1, Gemini-3-Pro-Preview, and Qwen3-Max, suggesting that stronger base models and proprietary training data still provide advantages under this challenging hard-case distribution.
General QA vs. Multi-Hop QA. CW-GRPO demonstrates consistent improvements on both General QA and Multi-Hop QA tasks compared to outcome-supervised and process-supervised baselines. On Multi-Hop QA, CW-GRPO yields the most pronounced gains. On both Qwen3-1.7B and Qwen3-8B, it improves performance on 2WikiMultiHopQA, Musique, and Bamboogle simultaneously, indicating stronger long-horizon reasoning and evidence aggregation. On General QA, CW-GRPO remains competitive across datasets. While it achieves the best overall General QA performance on Qwen3-8B, its improvement on Qwen3-1.7B is smaller than that of MT-PPO. This suggests that for simpler single-hop questions, dense process supervision can still be effective for very small models, whereas CW-GRPO shows clearer advantages as reasoning depth increases.
Summary. Overall, these results demonstrate that CW-GRPO provides a robust and scalable improvement over existing outcome- and process-supervised search agents, particularly for multi-hop reasoning under strict retrieval constraints. The consistent gains across model sizes highlight the effectiveness of contribution-weighted optimization in guiding search behavior, even in extremely challenging evaluation settings.